# The Wavefront Set: Bounds for the Langlands parameter

### v2

@@ 4.2. Symplectic groups
${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)=(2k_{P_{1}},4k_{P_{1}},\dots,2m_{P_{1}}^{(1)}k_{P_{1}})\sqcup(2k_{P_{2}},4k_{P_{2}},\dots,2m_{P_{2}}^{(2)}k_{P_{2}})$
@@
$d^{\vee}({\mathbb{O}}^{\vee}_{\pi})=(2k_{P_{1}},4k_{P_{1}},\dots,2m_{P_{1}}^{(1)}k_{P_{1}})+(2k_{P_{2}},4k_{P_{2}},\dots,2m_{P_{2}}^{(2)}k_{P_{2}}).$
@@
Claim (1) is just a restatement of the above procedure by Lust-Stevens, listed here for convenience. Claim (2) follows from formula (3.1) by a direct calculation using $j$-induction. More precisely, ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ is the saturation (concatenation in terms of the partitions) of the two Kawanaka wavefront set nilpotent orbits on the two factors $Sp(2n_{i})$ of the maximal parahoric.
@@
$j_{W(B)\times{W(D)}}^{W_{n}}\left(((1,2,\dots,m_{+})\times(1,2,\dots,m_{+}))\boxtimes(1,2,\dots,m_{-})\times(1,2,\dots,m_{-}-1))\right),$
@@
$\left((1,2,\dots,m_{+})+(1,2,\dots,m_{-})\right)\times\left((1,2,\dots,m_{+})+(1,2,\dots,m_{-}-1)\right),$
where $+$ has the meaning from Definition 4.2. Applying the combinatorial algorithm for the Springer representation for type $C$, we find that the corresponding symplectic nilpotent orbit is
$[2(m_{+}+m_{-})]\sqcup\begin{cases}[2(m_{+}-m_{-})],&m_{+}\geq m_{-},\\ [2(m_{-}-m_{+}-1)],&m_{+}<m_{-}\end{cases},$
@@
Claim (3) follows by computing Spaltenstein duality $d^{\vee}({\mathbb{O}}^{\vee}_{\pi})$ similarly to the proof of the previous proposition. In the case when $n_{2}=0$, so $m_{+}^{(2)}=m_{-}^{(2)}=0$, we get equality $d^{\vee}({\mathbb{O}}^{\vee}_{\pi})={}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$. ∎
We can now consider the general case by putting together the two cases in Propositions 4.5 and 4.4.
@@
Let $\lambda_{\pm}^{\vee}$ be the orthogonal partition for the dual nilpotent orbit given by the Lust-Stevens algorithm for the polynomials $P$ with degree equal to $1$. That is, $\lambda_{\pm}^{\vee}$ has an expression like in Proposition 4.5(1). Let $\mu_{\pm}$ be the symplectic partition for the wavefront set again for the $\deg P>1$ part as in Proposition 4.5(2).
Similarly, let $\lambda^{\vee}_{\text{uni}}$ and $\mu_{\text{uni}}$ be the orthogonal and symplectic partitions that encode the dual nilpotent orbit and the wavefront set, respectively, in the case when all polynomials have degree strictly greater than $1$, i.e., in the setting of Proposition 4.4(1),(2).
@@
$d^{\vee}({\mathbb{O}}^{\vee}_{\pi})$ $=d^{\vee}(\lambda^{\vee}_{\pm})+(\lambda^{\vee}_{\text{uni}})^{t}\geq\mu_{\pm}+\mu_{\text{uni}},$
since by Propositions 4.5 and 4.4, we know that
@@
In the case when $n_{2}=0$ (hyperspecial), equality holds for both, therefore for the combination. ∎
@@ 4.4. Special orthogonal groups
$\sigma=\begin{cases}j_{W(B)\times W(B)}^{W(B)}\left(((1,2,\dots,m_{+})\times(1,2,\dots,m_{+}))\boxtimes(1,2,\dots,m_{-})\times(1,2,\dots,m_{-}))\right),&n^{an}=1,\\ j_{W(D)\times W(D)}^{W(D)}\left(((1,2,\dots,m_{+})\times(1,2,\dots,m_{+}-1))\boxtimes(1,2,\dots,m_{-})\times(1,2,\dots,m_{-}-1))\right),&n^{an}=0,2.\end{cases}$
@@
$\sigma=\begin{cases}((1,2,\dots,m_{+})+(1,2,\dots,m_{-}))\times((1,2,\dots,m_{+})+(1,2,\dots,m_{-})),&n^{an}=1,\\ ((1,2,\dots,m_{+})+(1,2,\dots,m_{-}))\times((1,2,\dots,m_{+}-1)+(1,2,\dots,m_{-}-1)),&n^{an}=0,2.\end{cases}$
@@
Then (2) follows. Claim (3) is verified in the same way as for the symplectic case.

### v3

@@ 4.3. Symplectic groups
$^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)=(2m_{P_{1}}^{(1)}k_{P_{1}},\dots,4k_{P_{1}},2k_{P_{1}})\sqcup(2m_{P_{2}}^{(2)}k_{P_{2}},\dots,4k_{P_{2}},2k_{P_{2}})$
@@
$d^{\vee}({\mathbb{O}}^{\vee}_{\pi})=(2m_{P_{1}}^{(1)}k_{P_{1}},\dots,4k_{P_{1}},2k_{P_{1}})+(2m_{P_{2}}^{(2)}k_{P_{2}},\dots,4k_{P_{2}},2k_{P_{2}}).$
@@
Claim (1) is just a restatement of the above procedure by Lust–Stevens, listed here for convenience. Claim (2) follows from formula (3.1) by a direct calculation using $j$-induction. More precisely, $^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ is the saturation (concatenation in terms of the partitions) of the two Kawanaka wavefront set nilpotent orbits on the two factors $Sp(2n_{i})$ of the maximal parahoric.
@@
$j_{W(B)\times{W(D)}}^{W_{n}}\left((m_{+},\dots,2,1)\times(m_{+},\dots,2,1))\boxtimes(m_{-},\dots,2,1)\times(m_{-}-1,\dots,2,1))\right),$
@@
$\left((m_{+},\dots,2,1)+(m_{-},\dots,2,1)\right)\times\left((m_{+},\dots,2,1)+(m_{-}-1,\dots,2,1)\right),$
where ++ has the meaning from Definition 4.2. Applying the combinatorial algorithm for the Springer representation for type $C$, we find that the corresponding symplectic nilpotent orbit is
$2(m_{+}+m_{-})]\sqcup\begin{cases}[2(m_{+}-m_{-})],&m_{+}\geq m_{-},\\ [2(m_{-}-m_{+}-1)],&m_{+}<m_{-}\end{cases},$
@@
Claim (3) follows by computing Spaltenstein duality $d^{\vee}({\mathbb{O}}^{\vee}_{\pi})$ using formulas (4.7) and (4.3). We begin with the hyperspecial case: $n_{2}=0$, and $m_{+}^{(2)}=m_{-}^{(2)}=0$, and explain how we get an equality $d^{\vee}({\mathbb{O}}^{\vee}_{\pi})={}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$.
For simplicity, denote $a=m_{+}^{(1)}$, $b=m_{-}^{(1)}$. Then the partition for $\mathbb{O}_{\pi}^{\vee}$ is
$\underline{p}=[2a+1]\sqcup[2a-1]\sqcup[2b-1]\sqcup[2b-1].$
Firstly, notice that
$(2a+1,2a-1,2a-3,\dots,3,1)^{t}=(a+1,a,a,a-1,a-1,\dots,2,2,1,1),$
an identity that we will use repeatedly in the proof. Next, using (4.3), we get
$\underline{p}^{t}$ $=(a+1,a,a,a-1,a-1,\dots,1,1)+(a,a-1,a-1,\dots,1,1)+(2b,2(b-1),2(b-1),\dots,4,4,2,2)$ (4.8)
$=(2a+1,2a-1,2a-1,\dots,3,3,1,1)+(2b,2(b-1),2(b-1),\dots,4,4,2,2)$
[…]
Now it is easy to see that removing a $1$ and applying the $C$-collapse, we get
$d^{\vee}(\mathbb{O}_{\pi}^{\vee})=((\underline{p}^{t})^{-})_{C}=\begin{cases}[2(a+b)]\sqcup[2(a-b)],&a\geq b,\\ [2(a+b)]\sqcup[2(b-a)-2],&a<b,\end{cases}$
and this is exactly the partition for $^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$.
It remains to prove the inequality $d^{\vee}(\mathbb{O}_{\pi}^{\vee})\geq{}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ for arbitrary $m_{\pm}^{(i)}$.
Denote $\underline{p}=[2(m_{+}^{(1)}+m_{+}^{(2)})+1]\sqcup[2|m_{+}^{(1)}-m_{+}^{(2)}|-1]\sqcup[2(m_{-}^{(1)}+m_{-}^{(2)})-1]\sqcup[2|m_{-}^{(1)}-m_{-}^{(2)}|-1].$ Using (4.3), we have
$\underline{p}^{t}$ $=(m_{+}^{(1)}+m_{+}^{(2)}+1,m_{+}^{(1)}+m_{+}^{(2)},m_{+}^{(1)}+m_{+}^{(2)},\dots,1,1)+(|m_{+}^{(1)}-m_{+}^{(2)}|,|m_{+}^{(1)}-m_{+}^{(2)}|-1,|m_{+}^{(1)}-m_{+}^{(2)}|-1,\dots,1,1)+$
$(m_{-}^{(1)}+m_{-}^{(2)},m_{-}^{(1)}+m_{-}^{(2)}-1,m_{-}^{(1)}+m_{-}^{(2)}-1,\dots,1,1)+(|m_{-}^{(1)}-m_{-}^{(2)}|,|m_{-}^{(1)}-m_{-}^{(2)}|-1,|m_{-}^{(1)}-m_{-}^{(2)}|-1,\dots,1,1).$
This is a sum of four partitions. Set $\gamma_{+}$ to be the sum of the first two, and $\gamma_{-}$ the sum of the last two. Then $\underline{p}^{t}=\gamma_{+}+\gamma_{-}$,
$\gamma_{+}$ $=(2M_{+}+1,2M_{+}-1,2M_{+}-1,2M_{+}-3,2M_{+}-3,\dots,1,1)\sqcup(2m_{+},2m_{+},2m_{+}-2,2m_{+}-2,\dots,2,2),$
$\gamma_{-}$ $=(2M_{-},2M_{-}-2,2M_{-}-2,2M_{-}-4,2M_{-}-4,\dots,2,2)\sqcup(2m_{-}-1,2m_{-}-1,2m_{-}-3,2m_{-}-3,\dots,1,1),$
where $M_{\pm}=\max\{m_{\pm}^{(1)},m_{\pm}^{(2)}\}$, and $m_{\pm}=\min\{m_{\pm}^{(1)},m_{\pm}^{(2)}\}.$ Denote the two partitions in $\gamma_{+}$ by $\gamma_{+}^{M}$ and $\gamma_{+}^{m}$, respectively, so $\gamma_{+}=\gamma_{+}^{M}\sqcup\gamma_{+}^{m}$, and similarly for $\gamma_{-}$.
We need to compare $((\gamma_{+}+\gamma_{-})^{-})_{C}$ with the partitions describing the geometric wavefront set. We will use inequality (4.4). Notice that for $\gamma_{+}+\gamma_{-}$, there are two ways to apply this inequality and we will choose the appropriate way depending on the relation between the $m_{\pm}^{(i)}$’s. More precisely, if
  1. (1)
    $M_{+}=m_{+}^{(i)}$ and $M_{-}=m_{-}^{(i)}$, then $\alpha_{1}=\gamma_{+}^{M}$ and $\alpha_{2}=\gamma_{-}^{M}$;
  2. (2)
    $M_{+}=m_{+}^{(i)}$ and $M_{-}=m_{-}^{(i^{\prime})}$, $i\neq i^{\prime}$, then $\alpha_{1}=\gamma_{+}^{M}$ and $\alpha_{2}=\gamma_{-}^{m}$.
To compute $(\alpha_{1}+\alpha_{2})$ and $(\beta_{1}+\beta_{2})$, one case, namely when $\alpha_{1}=\gamma_{+}^{M}$ and $\alpha_{2}=\gamma_{-}^{M}$ has already appeared in (4.8). The other possibilities are of the form:
  1. (1)
    $(2a+1,2a-1,2a-1,\dots,3,3,1,1)+(2b-1,2b-1,2b-3,2b-3,\dots,1,1)$
    $=\begin{cases}(2(a+b),2(a+b)-2,2(a+b)-4,\dots,2(a-b)+2,2(a-b)+1,&\\ \ \ \ \ \ \ \ \ \ \ 2(a-b)-1,2(a-b)-1,\dots,3,3,1,1),&\ \ \ a\geq b,\\ (2(a+b),2(a+b)-2,2(a+b)-4,\dots,2(b-a),2(b-a)-1,&\\ \ \ \ \ \ \ \ \ 2(b-a)-3,2(b-a)-3,\dots,3,3,1,1),&\ \ \ a<b;\end{cases}$ (4.9)
  2. (2)
    $(2a,2a-2,2a-2,2a-4,2a-4,\dots,2,2)+(2b,2b,2b-2,2b-2,\dots,2,2)$
    $=\begin{cases}(2(a+b),2(a+b)-2,2(a+b)-4,\dots,2(a-b),&\\ \ \ \ \ \ \ \ \ \ \ 2(a-b)-2,2(a-b)-2,\dots,4,4,2,2),&\ \ \ a\geq b+1,\\ (2(a+b),2(a+b)-2,2(a+b)-4,\dots,2(b-a)+2,&\\ \ \ \ \ \ \ \ \ 2(b-a),2(b-a),\dots,4,4,2,2),&\ \ \ a\leq b;\end{cases}$ (4.10)
  3. (3)
    $(2a,2a,2a-2,2a-2,\dots,2,2)+(2b-1,2b-1,2b-3,2b-3,\dots,1,1)$
    $=\begin{cases}(2(a+b)-1,2(a+b)-1,2(a+b)-5,2(a+b)-5,\dots,2(a-b)+3,2(a-b)+3,&\\ \ \ \ \ \ \ \ \ \ \ 2(a-b),2(a-b),\dots,4,4,2,2),&\ \ \ a\geq b,\\ (2(a+b)-1,2(a+b)-1,2(a+b)-5,2(a+b)-5\dots,2(b-a)+3,2(b-a)+3&\\ \ \ \ \ \ \ \ \ 2(b-a)-1,2(b-a)-1,\dots,3,3,1,1),&\ \ \ a<b;\end{cases}$ (4.11)
Suppose $M_{+}=m_{+}^{(1)}$ and $M_{-}=m_{-}^{(1)}$. Then
$\gamma_{+}+\gamma_{-}\geq(\gamma_{+}^{M}+\gamma_{-}^{M})\sqcup(\gamma_{+}^{m}+\gamma_{-}^{m})$
The right hand side equals the union of (4.8), in which $a=m_{+}^{(1)}$ and $b=m_{-}^{(1)}$, and (4.11), in which $a=m_{+}^{(2)}$ and $b=m_{-}^{(2)}$. It is not difficult, albeit a bit tedious, to compute the $C$-collapse and find that
$((\gamma_{+}^{M}+\gamma_{-}^{M})\sqcup(\gamma_{+}^{m}+\gamma_{-}^{m}))^{-}_{C}$
equals exactly the disjoint union of the four partitions in the description of $^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$.
Notice that one could be tempted to use the inequality
$((\gamma_{+}^{M}+\gamma_{-}^{M})\sqcup(\gamma_{+}^{m}+\gamma_{-}^{m}))^{-}_{C}\geq(\gamma_{+}^{M}+\gamma_{-}^{M})^{-}_{C}\sqcup(\gamma_{+}^{m}+\gamma_{-}^{m}),$
noting that in this case $(\gamma_{+}^{m}+\gamma_{-}^{m})$ is already a partition of type $C$. Moreover, $(\gamma_{+}^{M}+\gamma_{-}^{M})^{-}_{C}$ equals, by (4.8), exactly the union of two partitions in the wavefront set of the first factor ($i=1$). But $\gamma_{+}^{m}+\gamma_{-}^{m}$ is smaller than the union of the two partitions in the wavefront set of the second factor ($i=2$). So the $C$-collapse can’t be separated in this computation.
Example 4.7.
If $m_{+}^{(1)}=8$, $m_{-}^{(1)}=5$, $m_{+}^{(2)}=6$, $m_{-}^{(2)}=2$, then $M_{+}=8$, $M_{-}=5$, $m_{+}=6$, $m_{-}=3$, and
$\gamma_{+}^{M}+\gamma_{-}^{M}$ $=(27,23,23,19,19,15,15,11,11,7,7,5,5,3,3,1,1)$
$\gamma_{-}^{m}+\gamma_{-}^{m}$ $=(17,17,13,13,9,9,6,6,4,4,2,2),$
so $(\gamma_{+}^{M}+\gamma_{-}^{M})\sqcup(\gamma_{-}^{m}+\gamma_{-}^{m})$ equals
(27,23,23,19,19,17,17,15,15,13,13,11,11,9,9,7,7,6,6,5,5,4,4,3,3,2,2,1,1).(27,23,23,19,19,17,17,15,15,13,13,11,11,9,9,7,7,6,6,5,5,4,4,3,3,2,2,1,1).
Removing a $1$ and applying the $C$-collapse, we get
$(26,$ $24,22,20,18,18,16,16,14,14,12,12,10,10,8,8,6,6,6,6,4,4,4,4,2,2,2,2)$
$=[26]\sqcup[18]\sqcup[6]\sqcup[6]$
$=[2(m_{+}^{(1)}+m_{-}^{(1)})]\sqcup[2(m_{+}^{(2)}+m_{-}^{(2)})]\sqcup[2(m_{+}^{(1)}-m_{-}^{(1)})]\sqcup[2(m_{+}^{(2)}-m_{-}^{(2)})],$
as required.
Suppose $M_{+}=m_{+}^{(1)}$ and $M_{-}=m_{-}^{(2)}$. Then
$\gamma_{+}+\gamma_{-}\geq(\gamma_{+}^{M}+\gamma_{-}^{m})\sqcup(\gamma_{+}^{m}+\gamma_{-}^{M})$
The right hand side equals the union of (4.9), in which $a=m_{+}^{(1)}$ and $b=m_{-}^{(1)}$, and (4.10), in which $a=m_{-}^{(2)}$ and $b=m_{+}^{(2)}$. Again we compute the $C$-collapse and find that
$((\gamma_{+}^{M}+\gamma_{-}^{M})\sqcup(\gamma_{+}^{m}+\gamma_{-}^{m}))^{-}_{C}$
equals the disjoint union of the four partitions in the description of $^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$. This completes the proof.
Example 4.8.
If $m_{+}^{(1)}=8$, $m_{-}^{(1)}=3$, $m_{+}^{(2)}=6$, $m_{-}^{(2)}=5$, then again $M_{+}=8$, $M_{-}=5$, $m_{+}=6$, $m_{-}=3$, but
$\gamma_{+}^{M}+\gamma_{-}^{m}$ $=(22,20,18,16,14,12,11,9,9,7,7,5,5,3,3,1,1)$
$\gamma_{-}^{M}+\gamma_{+}^{m}$ $=(22,20,18,16,14,12,10,8,6,4,2,2),$
so $(\gamma_{+}^{M}+\gamma_{-}^{M})\sqcup(\gamma_{-}^{m}+\gamma_{-}^{m})$ equals
(22,22,20,20,18,18,16,16,14,14,12,12,11,10,9,9,8,7,7,6,5,5,4,3,3,2,2,1,1).(22,22,20,20,18,18,16,16,14,14,12,12,11,10,9,9,8,7,7,6,5,5,4,3,3,2,2,1,1).
Removing a $1$ and applying the $C$-collapse, we get
$(22,$ $22,20,20,18,18,16,16,14,14,12,12,10,10,10,8,8,8,6,6,6,4,4,4,2,2,2,2)$
$=[22]\sqcup[22]\sqcup[10]\sqcup[2]$
$=[2(m_{+}^{(1)}+m_{-}^{(1)})]\sqcup[2(m_{+}^{(2)}+m_{-}^{(2)})]\sqcup[2(m_{+}^{(1)}-m_{-}^{(1)})]\sqcup[2(m_{+}^{(2)}-m_{-}^{(2)})],$
as required.
We can now consider the general case by putting together the two cases in Propositions 4.6 and 4.5.
@@
Let $\lambda_{\pm}^{\vee}$ be the orthogonal partition for the dual nilpotent orbit given by the Lust–Stevens algorithm for the polynomials $P$ with degree equal to $1$. That is, $\lambda_{\pm}^{\vee}$ has an expression like in Proposition 4.6(1). Let $\mu_{\pm}$ be the symplectic partition for the wavefront set again for the $\deg P=1$ part as in Proposition 4.6(2).
Similarly, let $\lambda^{\vee}_{\text{uni}}$ and $\mu_{\text{uni}}$ be the orthogonal (with an entry $1$ removed) and symplectic partitions that encode the dual nilpotent orbit and the wavefront set, respectively, in the case when all polynomials have degree strictly greater than $1$, i.e., in the setting of Proposition 4.5(1),(2).
@@
$d^{\vee}({\mathbb{O}}^{\vee}_{\pi})$ $=d^{\vee}(\lambda^{\vee}_{\pm})+(\lambda^{\vee}_{\text{uni}})^{t}\geq\mu_{\pm}+\mu_{\text{uni}},$ (4.12)
since by Propositions 4.6 and 4.5, we know that
@@
In the case when $n_{2}=0$ (hyperspecial), equality holds for both, therefore for the combination.
To justify the first equality in (4.12), write $\lambda^{\vee}_{\text{uni}}=\underline{p}_{A}\sqcup\underline{p}_{A}$, where $\underline{p}_{A}=\bigsqcup_{P}[m_{P}^{(1)}+m_{P}^{(2)}]^{k_{P}}\sqcup[|m_{P}^{(1)}-m_{P}^{(2)}|-1]^{k_{P}}$ in the notation of Proposition 4.5. Then we may regard $\mathbb{O}_{\pi}^{\vee}$ as the $G^{\vee}$-saturation of an orbit $\mathbb{O}^{\vee}_{0}$ on a maximal Levi subgroup of type $GL(n-m)\times SO(2m+1)$, where the nilpotent $GL$-part of $\mathbb{O}_{0}^{\vee}$ is given by $\underline{p}_{A}$ and the $SO$-part by $\lambda_{\pm}^{\vee}$. Using Lemma 2.1(3), it follows that
$d^{\vee}(\mathbb{O}^{\vee}_{\pi})$ $=\operatorname{ind}_{GL(n-m)\times Sp(2m)}^{Sp(2n)}(\underline{p}_{A}^{t},d^{\vee}(\lambda^{\vee}_{\pm}))=(\underline{p}_{A}^{t}+\underline{p}_{A}^{t}+d^{\vee}(\lambda^{\vee}_{\pm})_{C}$ (4.13)
$=((\lambda_{\text{uni}}^{\vee})^{t}+d^{\vee}(\lambda^{\vee}_{\pm}))_{C}=(\lambda_{\text{uni}}^{\vee})^{t}+d^{\vee}(\lambda^{\vee}_{\pm}),$
where the last equality follows because all the parts in both partitions are even. ∎
@@ 4.5. Special orthogonal groups
$\sigma=\begin{cases}j_{W(B)\times W(B)}^{W(B)}\left(((m_{+},\dots,2,1)\times(m_{+},\dots,2,1))\boxtimes(m_{-},\dots,2,1)\times(m_{-},\dots,2,1))\right),&n^{an}=1,\\ j_{W(D)\times W(D)}^{W(D)}\left(((m_{+},\dots,2,1)\times(m_{+}-1,\dots,2,1))\boxtimes(m_{-},\dots,2,1)\times(m_{-}-1,\dots,2,1))\right),&n^{an}=0,2.\end{cases}$
@@
$\sigma=\begin{cases}((m_{+},\dots,2,1)+(m_{-},\dots,2,1))\times((m_{+},\dots,2,1)+(m_{-},\dots,2,1)),&n^{an}=1,\\ ((m_{+},\dots,2,1)+(m_{-},\dots,2,1))\times((m_{+}-1,\dots,2,1)+(m_{-}-1,\dots,2,1)),&n^{an}=0,2.\end{cases}$
@@
Then (2) follows. Claim (3) is verified by analogous computations to the symplectic case.